(1) A tank is filled with 300 liters (L) of contaminated water initially containing 3 kg of toxin. Pure water is pumped in to the tank at a rate of 40 L/min, mixes instantaneously, and the mixture is pumped out at the same rate (40 L/min). Let y(t) be the amount of toxin present in the tank at time t. Write the differential equation and the initial value that models this scenario.

(a) Find the particular solution, y(t). Put a box around your answer.
Answer #1

roupa hie - (78) = (a) hif [9.0=nle so (esh zania (trad = (4) h t he = s-ah & Io ho -] = (wh-cə Red